OpenStudy (anonymous):

What is the larger solution to the equation I6-x/6I=11

jhonyy9 (jhonyy9):

so this is (16-x)/61 =11 ?

jhonyy9 (jhonyy9):

right ?

jhonyy9 (jhonyy9):

so if yes is right than will be 16-x=11*61 x=16-11*61 x=16-671 =- 655 x=- 655

OpenStudy (anonymous):

it looks like an absolute value function. That means that 6-(x/6) = ±11 let's start with +11 : 6-(x/6) = 11 -x/6 = 5 x = -30 if -11, 6-(x/6) = -11 -x/6 = -17 x = 102 if your equation was |(6-x)/6| = 11, the answers will be different tho (brackets are important here)

jhonyy9 (jhonyy9):

so if this is modul of ---- - so than (6-x)/6 =+/-11 - so 6-x=66 --- x=-60 - so 6-x=-66 --- x= 72